OpenStudy (anonymous):

Review question: A 48g bullet is fired with a speed of 500 m/s into a 3.6 kg sandbag pendulum that is free to swing. To what maximum vertical height will the endulum rise? Assume the bullet remains mbedded in the sandbag and that the rope is long.

OpenStudy (shamim):

kinetic energy of billet is (1/2)mv^2=0.5*48*10^-3*500^2

OpenStudy (shamim):

this kinetic energy of a bullete will b equal to the potentisl energy of sandbag

OpenStudy (anonymous):

yeah i did that, but i did not get the right answer. the answer is 2.2 m

OpenStudy (shamim):

ok

OpenStudy (anonymous):

so how to get the right answeR?

OpenStudy (shamim):

i m trying

OpenStudy (shamim):

now potential energy of sandbag with bullete is

OpenStudy (shamim):

mgh=(3.6+48*10^-3)*9.8*h

OpenStudy (shamim):

did u c the value of m

OpenStudy (anonymous):

Which of the following is not used to hold liquids?

OpenStudy (anonymous):

Which of the following is not used to hold liquids?

OpenStudy (anonymous):

a. Scientific models can save money and lives. b. Scientific models allow scientists to test their predictions. c. Scientific models allow scientists to study systems that no longer exist. d. Scientific models are used rarely.

OpenStudy (shamim):

u should post ur question in home page